Which of the following are partitions of , the set of real numbers? Explain your answers.
a. {In : n ∈ ℤ}, where In = {x ∈ ℝ : n ≤ x ≤ n + 1}
b. {Jn : n ∈ ℤ }, where Jn = {x ∈ ℝ: n ≤ x < n + 1}
1
Expert's answer
2020-11-22T18:44:40-0500

By defenition, a partition of a set "\mathbb R" is a set of non-empty subsets of "\mathbb R"  such that every element "x"  in "\mathbb R" is in exactly one of these subsets.


a. Let "\{I_n : n \in\mathbb Z\}", where "I_n = \{x \in\mathbb R : n ≤ x ≤ n + 1\}". It is not a partition of a set "\mathbb R". Indeed, "I_0=[0,1]" is a unique set that contains the real number "0.5" and "I_1=[1,2]" is a unique set that contains the real number "1.5". On the other hand, a real number 1 belong to both set "I_0" and "I_1."


b. Let "\{J_n : n \in\mathbb Z\}", where "J_n = \{x \in\mathbb R : n ≤ x < n + 1\}". Then it is a partition of a set "\mathbb R". Let "x" be a real number.The floor function "\lfloor x \rfloor" is defined to be the greatest integer less than or equal to the real number "x". Let "n=\lfloor x \rfloor". Then "x\in[n,n+1)". Since "[n,n+1)" and "[m,m+1)" are disjoint for different integers "n" and "m", "\{J_n : n \in\mathbb Z\}" is a partition.